How do I solve T=2pi(sqrt(L/G)) ?

Answer by stanbon(75887)

solve T=2pi(sqrt(L/G))
Divide both sides by 2pi
sqrt(L/G) = T/(2pi)
Square both sides to get:
L/G = [T/(2pi)]^2
---
solve for "L":
L = G[T/2pi]^2
=======================
Cheers,
Stan H.